# Training Convolutional Neural Networks with the Forward-Forward Algorithm

### 25 Dec 2023 v2

@@ 2.3 Two Ways of Inference
There are two ways how an FF trained CNN or DNN can be used for inference: the linear classifier and the goodness evaluation. In the first case, the $H$ neurons of every layer except the first are fully-connected with an output layer of $N$ nodes, equal to the number of labels. The connecting weights, $H$ times $N$ in number, are trained by evaluating the neuron activations using a cross-entropy loss. If not mentioned differently, this is the default method of inference used in this paper.
For inference with goodness evaluation, each image is exposed $N$ times to the neural network, each time superimposed with another of the $N$ possible labels. The goodness parameter is computed for each label and the image is associated with the label characterized by the highest goodness value:
$\text{label}_{\text{\,correct}}=\text{argmax}\left(\begin{bmatrix}f_{0}\\ f_{1}\\ \vdots\\ f_{9}\end{bmatrix}\right),$
where the goodness is expressed as $f_{m}=\sum^{H}_{i=1}y_{i}^{2}$ for each associated label $m$ and $H$ is again the number of all neurons, except those from the first layer.
@@ 2.4 Hardware and Software
The code for our FF trained CNN is implemented in Python using the PyTorch library (pytorch). The source code of Loewe (2023) for an FF trained, fully connected DNN was used as the starting point and our code is available on GitHub [^4] Code will be available after publication in a peer-reviewed journal.. All analysis presented here was performed on a desktop workstation with an AMD Ryzen 9 5900X 12-Core Processor with 128 GB RAM, and an NVIDIA RTX 3080 GPU with 12 GB RAM.
@@ 3 Results
We first report the configuration which achieved the highest accuracy, then the search through the space of hyperparameters (using only our validation data set) leading to the optimal configuration. We close by demonstrating the ability of FF trained CNN to implement Class Activation Maps, which is a method from the explainable AI toolbox.
@@ 3.1 Performance of the Optimized Configuration
The hyperparameters optimization leads to the following configuration for the FF trained CNN: three convolution layers of each 128 filters with a kernel dimension of 7x7 pixels. After training for 200 epochs with a batch size of 50, Adam optimizer with a learning rate = $5\times 10^{-5}$, and the label set 1 (Figure 2) with intensity $K$ = 35%, we obtain 99.20% accuracy for the validation data set and 99.16% for the test data set using the goodness approach for inference. While having a significantly shorter run time, inference with the linear classifier approach provides slightly worse results, achieving accuracy values of 99.14% and 99.00% for validation and test data sets, respectively. For comparison, using again a three layer CNN of constant size but trained with BP we obtain a validation accuracy of 99.13%. Here the search for optimal hyperparameters resulted in 16 filters of 5x5 pixels, Adam optimizer with learning rate = $10^{-3}$, 200 epochs with batch size = 50.
Figure 3 provides a more detailed picture of this comparison. Figure 3 A shows that the performance of FF training increases monotonously with the number of filters per layer while BP training accuracy decreases slightly under the same conditions. The latter is most likely due to increasing overfitting; its absence indicates that FF training might either be more robust against overfitting, or that it makes less efficient use of its number of trainable parameters.
The confusion matrix in Figure 3 B provides insight into the classification performance, revealing that labels 5 and 9 exhibit the least accurate classifications (lower than 98.20%), while labels 6, 7 and 1 are characterized by the highest accuracy levels (higher than 99.50%).
@@
After 200 epochs, the discrimination loss of the FF trained network reaches a plateau for all considered layers (Figure 3 C). This convergence of the training is confirmed by a training run with 750 epochs which results in no substantial changes in accuracy. Moreover, he accuracy value on the training data reaches close to 100% (depicted in Figure 3 D with a green line) for 200 epochs. The discrimination accuracy values of the layers (red and blue lines in Figure 3 D) corroborate this result. They also hint at a slightly different dynamics of the two layers 2 and 3 with layer 2 initially learning faster, but layer 3 achieving a higher faculty of discrimination on the long run.
@@
Figure 3: The best MNIST performance of an FF-trained CNN architecture is comparable to the results of a backpropagation trained CNN of the same architecture. (A) shows the accuracy values obtained for CNN with three convolutional layers as a function of the number of filters in each layer after being trained for 200 epochs with batch size 50. Filter size is 7x7, the learning rate is set to the respective optimal value of $5\times 10^{-5}$ for FF and $10^{-3}$ for BP. FF trained networks use label set 1 and a label intensity $K$ of 35%. The values reported for BP and FF using the linear classifier are gathered from the validation data. The green data point shows the same network but with inference using the goodness comparison. Its 99.16% accuracy ws measured using the test data. (B) reports the confusion matrix of the FF-trained, optimal configuration CNN using the linear classifier for inference (C) shows the loss computed for the discrimination between positive and negative training data for each hidden layer contributing to the training (red and blue lines), and the combined loss used during training (green line). (D) displays the discrimination accuracy of the same hidden layers (red and blue lines), and the total accuracy obtained during training (green line).
@@ 3.2 Influence of Hyperparameters
The performance of BP trained neural networks depends on the value of the hyperparameters used during training. There is no reason to expect this to be different for FF training. This Section discusses the search we performed in order to obtain the optimal result reported in Section 3.1.
This search also included the exploration of some architectural options which are not hyperparameters in the strict sense. We limited our search to networks of three layers of convolutional filters, all with the same number of filters per layer. Moreover, we did not include max pooling layers because the tests described in Appendix C indicate that they decrease accuracy.
As optimizer we tested stochastic gradient descent and Adam with the latter consistently outperforming the former. For the learning rate of Adam we evaluated the range between $10^{-2}$ and $10^{-7}$. A learning rate of $5\times 10^{-5}$ yielded the best results and was therefore chosen as default. As in Section 3.1, we found that after training for 200 epochs the network had converged in all configurations tested, which made this value also our default.
@@ 3.2.1 Filter Dimensions, Layer Width, and Batch Size
First, we describe the interaction of the two architectural parameters filter dimension and number of neurons in each layer with the training parameter batch size. Figure 4 A shows that the accuracy decreases with decreasing filter size. A possible explanation is that the smaller filters have more problems identifying the wavelengths of the label waves. For 7x7 and 5x5 filters we additionally observe an increase in accuracy for smaller batch sizes. We speculate that this might be a generic feature due to the dual nature of the training data.
Figure 4 B confirms for all filter sizes the increase of accuracy with the number of filters per layer, which was first shown in Figure 3 A. Again the accuracy decreases with decreasing filter size.
Figure 4: FF-trained CNN benefit from small batch sizes and a large number of filters. (A) Accuracy as a function of batch size and filter size (3x3, 5x5, or 7x7 pixels) in a network of three layers of each 64 filters. (B) Accuracy values obtained by varying the number of filters per layer (16, 32, 64, 128) and the filter size using a batch size of 50.
@@
As described in Section 2.2 there are too ways for making inference with a FF trained neural network. Table 1 shows that the goodness computation approach beats the linear classifier over a variety of different CNN configurations. This finding agrees with the result of Brenig and Timofte (2023).
However, the slightly superior results go together with a ten times larger computational cost of the goodness comparison because each image in the data set must be processed for each label in order to perform the classification task. Consequentially, the goodness approach might only be justified for classification task with a limited number of labels.
@@
16 filter of size 5x5, batch size 50 98.16 98.75
64 filter of size 3x3, batch size 50 98.30 98.82
64 filter of size 5x5, batch size 200 98.67 98.94
Table 1: Using the goodness approach yields a higher accuracy during inference than the linear classifier, albeit at 10 times higher computational cost.
@@ 3.2.3 Effect of the Relative Labeling Intensity K
In studying the effect of the relative label intensity $K$, we focus on the hyperparameters that exhibit the highest accuracy values: 64 or 128 convolutional filters per layer, filter size 7x7, and batch sizes of 25 and 50 images. The results presented in Figure 5 show that optimizing $K$ leads to an increase in accuracy of up to 0.2%. The optimal value of $K$ does depend on layer width but not on batch size. Specifically, the 64x64x64 architectures achieves the highest accuracy for $K\approx 65\%$, the 128x128x128 networks for $K\approx 35\%$. This relatively large difference was unexpected.
@@ 3.3 Class Activation Maps
Class Activation Maps (CAMs) are visual representations that highlight the regions of an input image which contribute most to the prediction of a given label. CAMs are obtained by summing up the feature maps generated by the convolutional layers, each weighted with the corresponding weights associated with a specific label. The idea is that each feature map decodes specific spatial characteristics of the input image and the weights quantify how much these characteristics contributes to the recognition of the target class.
In BP trained CNN, CAMs are typically obtained by applying a global average pooling layer after the last convolutional layer and then using the weights connecting this layer to a softmax activation output layer. Here we train a CNN with FF and a linear classifier for inference. The weights of the linear classifier connecting the individual pixels in our feature map with the set of ten output neurons are exactly the weights we need to asses the role of the corresponding pixel for a given prediction.
Figure 6 provides four examples of CAMs of correctly identified images. For the digit 1, (Figure 6 A and B) the entire vertical shape contributes to the correct inference. In contrast, for the digit 2 in Figure 6 C three distinct areas (the upper, the bottom-left and the bottom right parts of the number) contribute to the correct labeling. Similarly expressive regions can be identified for the digits 7 and 9 in Figures 6 E and G.
Figure 6: Class activation maps (CAMs) of a FF trained CNN show which image regions are considered beneficial (yellow) or deleterious (pink) by the network for making its prediction. (A), (C) (E), and (G) display four input images. (B), (D) (F), and (H) are the corresponding CAMs. All examples are from a network with 16 convolutional neurons per layer, filter size 5x5, and which was trained with a batch size of 50.
CAMs also show that the different layers of the FF-trained CNN provide similar but distinct information for the classification task. For instance, when considering the digit 7 in Figure 7, the second layer of the network provides more information on the inner portion of the horizontal line, while the third layer responds more to the boundaries of that horizontal line.
Figure 7: Class Activation Maps show that the different layer of the FF-trained CNN provide similar, but yet distinguishable information. (A) shows the CAM obtained from considering both layer 2 and layer 3 together. (B) and (C) show the CAMs obtained from layer 2, respectively layer 3 alone.
@@ 4 Discussion and Conclusions
CNNs are considered the gold standard in deep learning-based image analysis. For instance, in biomedical imaging, they overcome the drawbacks of subjective analysis in the semi-quantitative visual inspection of samples (Gurcan et al., 2009) and they support experts during their daily clinical routine by reducing their workload (Shmatko et al., 2022). Furthermore, their exploitation of the spatial information within images makes them suitable for the deployment of explainable AI tools (such as Class Activation Maps) which highlight the image regions contributing most significantly to the classification outcome.
Our implementation of FF trained CNN shows that with the right choice of hyperparameters, this technique is competitive with backpropagation. These results were obtained without implementing all the possible and suggested optimizations such as enforcing symmetry of the loss function (Lee and Song, 2023) or choosing hard, i.e. easily confused, labels for the negative data set as suggested by Hinton (2022). We propose that our work shows the potential of FF trained CNNs to address real world computer vision problems. If this technique will in some applications supersede BP is an open question. We believe that this potential exists, especially in the cases of neuromorphic hardware and unsupervised learning.
A better understanding of the FF training will however also expand our understanding of the generic concept of neuronal information processing in all its breadth from biological systems to reservoir computing. The here demonstrated capability of implementing Class Activation Maps provides a first window into these research topics. Achieving deeper insights will also mean to understand how the two innovations of FF, providing positive and negative labels and computing a locally defined goodness parameter, contribute to its success individually and synergetically (Tosato et al., 2023). Moreover, a better understanding why it is beneficial to exclude the first layer during the goodness computation (c.f. Appendix B) would be desirable.
Subsequent work on FF training should also address its ability to train deeper networks, most likely expanding on the work of Lorberbom et al. (2023). Also the ability of FF training to work with larger and more complex data sets needs to be explored. Finally, its connection to biological neuronal systems (Ororbia and Mali, 2023; Ororbia, 2023) seems a promising research direction.
@@
This project has received funding from the Innovative Medicines Initiative 2 Joint Undertaking (JU) under grant agreement number 101034427. The JU receives support from the European Union’s Horizon 2020 research and innovation program and EFPIA. The JU is not participating as a contracting authority in this procurement. This project was also funded by the Ministry for Science and Culture of Lower Saxony as part of the project “Agile, bio-inspired architectures” (ABA).
@@ Appendix A A Second Way of Encoding the Labels
Figure 8: Label set 2 superimposed on an image of the digit 7. The 10 different labels share the same wavelength and differ only in their orientation. Only the label 7 is part of the positive data set, while one of the other images is randomly selected for the negative data set.
@@
128 filters of size 5x5, batch size 15 98.98 98.95
64 filters of size 5x5, batch size 15 98.83 98.81
32 filters of size 3x3, batch size 50 97.95 97.94
32 filters of size 5x5, batch size 50 98.50 98.47
32 filters of size 5x5, batch size 100 98.40 98.40
Table 2: FF-trained CNNs achieve higher accuracies when labeled with set 1 instead of set 2. Five different sets of hyperparameters were trained, the best results is highlighted.

### 07 Jan 2024 v3

@@ 2.3 Two ways of inference
There are two ways how a FF trained CNN or DNN can be used for inference: the linear classifier and the goodness evaluation. In the first case, $H$ neurons of every layer (except the first) are fully-connected with an output layer of $N$ nodes, equal to the number of labels. The connecting weights, $H$ times $N$ in number, are trained by evaluating the neuron activations using a cross-entropy loss. This is the default method of inference used in this paper, unless mentioned otherwise. For inference with goodness evaluation, each image is exposed $N$ times to the neural network, each time superimposed with another of the $N$ possible labels, and the goodness parameter is computed for each label $m$. The image is then associated with ${label}_{\,correct}$, which is the label characterized by the highest goodness value, and is defined as:
${label}_{\,correct}=\text{argmax}\left(\begin{bmatrix}f_{0}\\ f_{1}\\ \vdots\\ f_{8}\\ f_{9}\end{bmatrix}\right),$
where for each associated label $m$, the goodness is expressed as $f_{m}=\sum^{H}_{i=1}y_{i}^{2}$, and $H$ is the number of all neurons, except those from the first layer.
@@ 2.4 Hardware and software
The code for our FF trained CNN is implemented in Python using the PyTorch library (pytorch). The source code of Loewe (2023) for an FF trained, fully connected DNN was used as the starting point, and our code is available on GitHub [^4] Code will be available after publication in a peer-reviewed journal.. All analysis presented here was performed on a desktop workstation with an AMD Ryzen 9 5900X 12-Core Processor with 128 GB RAM, and an NVIDIA GeForce RTX 3080 GPU with dedicated 12 GB RAM.
@@ 3 Results
We first report the configuration which achieved the highest accuracy, followed by the search through the space of hyperparameters (using only our validation data set) leading to the optimal configuration. We close by demonstrating the ability of FF trained CNNs to implement Class Activation Maps, which is a method from the explainable AI toolbox.
@@ 3.1 Performance of the optimized configuration
The hyperparameter optimization leads to the following configuration for the FF trained CNN: three convolution layers of each 128 filters with a kernel dimension of 7x7 pixels. After training for 200 epochs with a batch size of 50 using the Adam optimizer with a learning rate of $5\times 10^{-5}$, and the label set 1 (Figure 2) with intensity $K$ of 35%, we obtain 99.20% accuracy for the validation data set and 99.16% for the test data set using the goodness approach for inference. While having a significantly shorter run time, inference with the linear classifier approach provides slightly worse results, achieving accuracy values of 99.14% and 99.00% for validation and test data sets, respectively. For comparison, using again a three layer CNN of constant size, but trained with BP, we obtain a validation accuracy of 99.13%. Here the search for optimal hyperparameters resulted in 16 filters of 5x5 pixels, Adam optimizer with learning rate of $10^{-3}$, 200 epochs with batch size of 50.
@@
Figure 3 provides a more detailed picture of this comparison. Figure 3A shows that the performance of FF training increases monotonously with the number of filters per layer, while BP training accuracy decreases slightly under the same conditions. The latter is most likely due to increasing overfitting; its absence indicates that FF training might either be more robust against overfitting, or that it makes less efficient use of its number of trainable parameters. The confusion matrix in Figure 3B provides insight into the classification performance, revealing that labels 4 and 9 exhibit the least accurate classifications (lower than 98.70%), while labels 1, 3 and 7 are characterized by the highest accuracy levels (higher than 99.20%). After 200 epochs, the discrimination loss of the FF trained network reaches a plateau for all considered layers (Figure 3C). This convergence of the training is confirmed by a training run with 750 epochs which results in no further substantial changes in accuracy. Moreover, the accuracy value on the training data reaches close to 100% (depicted in Figure 3D with a green line) for 200 epochs. The discrimination accuracy values of the layers (red and blue lines in Figure 3D) corroborate this result. They also hint at a slightly different interplay between the dynamics of layers 2 and 3, with layer 2 initially learning faster, but layer 3 achieving a higher faculty of discrimination on the long run.
@@
Figure 3: The best MNIST performance of an FF-trained CNN architecture is comparable to the results of a backpropagation trained CNN of the same architecture. (A) shows the accuracy values obtained for CNN with three convolutional layers as a function of the number of filters in each layer, after being trained for 200 epochs with batch size 50. Filter size is 7x7, the learning rate is set to the respective optimal value of $5\times 10^{-5}$ for FF and $10^{-3}$ for BP. FF trained networks use labels from set 1 and a label intensity $K$ of 35% (our defaults). The values reported for BP and FF are gathered from the validation data. The green data points shows the results related to the FF trained network, with inference using the goodness comparison. In this scenario, 99.16% accuracy was achieved with 128 filters per layer using the test data as shown by the corresponding confusion matrix reported in (B). (C) shows the loss computed for the discrimination between positive and negative training data for each hidden layer contributing to the training (red and blue lines), and the combined loss used during training (green line). (D) displays the discrimination accuracy of the same hidden layers (red and blue lines), and the total accuracy obtained during training (green line).
@@ 3.2 Influence of hyperparameters
This section discusses the search for optimal hyperparameters we performed in order to obtain the result reported in Section 3.1. The performance of BP trained neural networks depends on the specific value of the hyperparameters used during training. There is no reason to expect this to be different for FF training. This search also included the exploration of some architectural options which are not hyperparameters in the strict sense. We limited our search to networks of three layers of convolutional filters, all with the same number of filters per layer. Moreover, we did not include max pooling layers, because the tests described in Appendix C indicate that they decrease accuracy. For the choice of an optimizer, we tested Stochastic Gradient Descent and Adam optimizer, with the latter consistently outperforming the former. For the learning rate of Adam optimizer, we evaluated the range between $10^{-2}$ and $10^{-7}$. A learning rate of $5\times 10^{-5}$ yielded the best results and was therefore chosen as default. As in Section 3.1, we found that after training for 200 epochs the network had converged in all configurations tested, which made this value also our default.
@@ 3.2.1 Filter dimensions, layer width, and batch size
First, we describe the interaction of the two architectural parameters, namely, filter dimension and number of neurons in each layer with the training parameter batch size. Figure 4A shows that the accuracy decreases with lower filter size. A possible explanation is that the smaller filters have more problems in identifying the wavelengths of the label waves. For 7x7 and 5x5 filters we additionally observe an increase in accuracy for smaller batch sizes. We speculate that this might be a generic feature due to the dual nature of the training data. Figure 4B confirms the increase of accuracy with the number of filters per layer for all filter sizes, which was first shown in Figure 3A, with the accuracy decreasing with smaller filter size.
Figure 4: FF-trained CNN benefits from small batch sizes and a large number of filters. (A) Accuracy as a function of batch size and filter size (3x3, 5x5, or 7x7 pixels) in a network of three layers of each 64 filters. (B) Accuracy values obtained by varying the number of filters per layer (16, 32, 64, 128) and the filter size using a batch size of 50.
@@
As described in Section 2.2 there are two ways for making inference with a FF trained neural network. Table 1 shows that the goodness computation approach beats the linear classifier over a variety of different CNN hyperparameters configurations, and this finding agrees with the proposal of Hinton (2022) and the result of Brenig and Timofte (2023). However, the slightly superior results go together with a ten times larger computational cost of the goodness comparison, because each image in the data set must be processed for each label in order to perform the classification task. Consequentially, the goodness approach might only be justified for classification task with a limited number of labels.
@@
128 filter of size 3x3, batch size 25 98.63 98.74
Table 1: Using the goodness approach yields a higher accuracy during inference than the linear classifier, albeit at 10 times higher computational cost. Three different sets of hyperparameters were trained, with the best results highlighted in bold.
@@ 3.2.3 Effect of the relative labeling intensity K
In studying the effect of the relative label intensity $K$, we focus on the hyperparameters that exhibit the highest accuracy values: 32, 64 or 128 convolutional filters per layer, filter size 7x7, and batch sizes of 25 and 50 images. The results presented in Figure 5 show that optimizing $K$ leads to an increase in accuracy of up to 0.2%. The optimal value of $K$ does depend on layer width but not on batch size. Specifically, the 32 neurons per layer architectures achieves the highest accuracy for $K$ approximately $20\%$, for 64 neurons per layer the optimal $K$ increases to approximately $65\%$, for 128 neurons per layer it drops again to approximately $35\%$. This non-monotonic behaviour does not lend itself to a simple explanation.
@@ 3.3 Class activation maps
Class Activation Maps (CAMs) are visual representations that highlight the regions of an input image which contribute most to the prediction of a given label. CAMs are obtained by summing up the feature maps generated by the convolutional layers, each weighted with the corresponding weights associated with a specific label. The main underlying idea is that each feature map decodes specific spatial characteristics of the input image, and as a result, the weights quantify how much these characteristics contributes to the recognition of the target class. In BP trained CNNs, CAMs are typically obtained by applying a global average pooling layer after the last convolutional layer, followed by using the weights connecting this layer to a Softmax activation output layer.
Here we train a CNN with FF and a linear classifier for inference. The weights of the linear classifier connecting the individual pixels in our feature map with the set of ten output neurons are exactly the weights we need to assess the role of the corresponding pixel for a given prediction. Figure 6 provides four examples of CAMs of correctly identified images. For the digit 1, (Figure 6A and Figure 6B) the entire vertical shape contributes to the correct inference. In contrast, for the digit 2 in Figure 6C three distinct areas (the upper, the bottom-left and the bottom right parts of the number) contribute to the correct labeling. Similarly expressive regions can be identified for the digits 7 and 9 in Figure 6E and Figure 6G.
Figure 6: Class activation maps (CAMs) of a FF trained CNN show which image regions are considered beneficial (yellow) or deleterious (pink) by the network for making its prediction. (A), (C), (E), and (G) display four input images. (B), (D), (F), and (H) are their corresponding CAMs. All examples are from a network with 16 convolutional neurons per layer, filter size 5x5, and trained with a batch size of 50.
CAMs also show that the different layers of the FF trained CNN provide similar, but distinct information for the classification task. For instance, when considering the digit 7 in Figure 7, the second layer of the network provides more information on the inner portion of the horizontal line, while the third layer responds more to the boundaries of that horizontal line.
Figure 7: Class activation maps show that the different layers of the FF-trained CNN provide similar, but yet distinguishable information. (A) shows the CAM obtained from considering both layer 2 and layer 3 together. (B) and (C) show the CAMs obtained respectively only from layer 2 and layer 3.
@@ 4 Discussion and conclusions
CNNs are considered the gold standard in deep learning-based image analysis. For instance, in biomedical imaging, they overcome the drawbacks of subjective analysis in the semi-quantitative visual inspection of samples (Gurcan et al., 2009), and they support experts during their daily clinical routine by reducing their workload (Shmatko et al., 2022). Furthermore, their exploitation of the spatial information within images makes them suitable for the deployment of explainable AI tools (such as class activation maps), which highlight the image regions contributing most significantly to the classification outcome. Our implementation of FF trained CNN shows that with the right choice of hyperparameters, this technique is competitive with backpropagation. These results were obtained without implementing all the possible and suggested optimizations such as enforcing symmetry of the loss function (Lee and Song, 2023) or choosing hard, i.e. easily confused, labels for the negative data set, as suggested by Hinton (2022). We propose that our work shows the potential of FF trained CNNs to address real world computer vision problems. An open question remains if this technique will supersede BP in specific applications. We believe that this potential exists, especially in the cases of neuromorphic hardware and unsupervised learning.
A better understanding of the FF training will however also expand our understanding of the generic concept of neuronal information processing in all its breadth from biological systems to reservoir computing. The demonstrated capability to implement class activation maps offers an initial insight into these research topics. Achieving deeper insights will also mean to understand how the two innovations of FF, providing positive and negative labels and computing a locally defined goodness parameter, contribute to its success individually and synergetically (Tosato et al., 2023). Moreover, a better understanding why it is beneficial to exclude the first layer during the goodness computation (c.f. Appendix B) would be desirable. Subsequent work on FF training should also address its ability to train deeper networks, most likely expanding on the work of Lorberbom et al. (2023). Also the ability of FF training to work with larger and more complex data sets needs to be explored. Finally, its connection to biological neuronal systems (Ororbia and Mali, 2023; Ororbia, 2023) seems a promising research direction.
@@
This project has received funding from the Innovative Medicines Initiative 2 Joint Undertaking (JU) under grant agreement number 101034427. The JU receives support from the European Union’s Horizon 2020 research and innovation program and EFPIA. The JU is not participating as a contracting authority in this procurement. This project was also funded by the Ministry for Science and Culture of Lower Saxony as part of the project “Agile, bio-inspired architectures” (ABA). The authors also thank Christian Dullin for insightful discussions.
@@ Appendix A A second way of encoding the labels
Figure 8: Label set 2 superimposed on an image of the digit 7. The 10 different labels shown here share the same wavelength and differ only in their orientation. Only the label 7 from the MNIST data set is a part of the positive data set, while one image from the 10 images shown is randomly selected for the negative data set.
@@
32 filters of size 3x3, batch size 50 98.34 98.36
32 filters of size 7x7, batch size 25 98.72 98.53
64 filters of size 3x3, batch size 50 98.30 98.22
128 filters of size 3x3, batch size 50 98.43 98.38
Table 2: FF trained CNNs achieve higher accuracy values when labeled with set 1 instead of set 2. Four different sets of hyperparameters were trained, with the best results highlighted in bold.